CAT Chat Minutes

August 16, 2004

 

General Information
Machine Stats
   Delivered beam:   1676 hrs
   Percentage of scheduled time:   98.53%
   Downtime during period:   25.07 hrs.
   Mean time between faults:   59.84 hrs.

Faults:

# 47   08/10 16:00   To   08/13 22:56    78.94  

  S10B:V2:PS Trip      [PS]

  1.14 Supply swapped, refill, resume top-up.

# 48   08/14 00:04   To   08/16 09:05    57.01  

  Fill Ongoing             

  -0.00  

Beamline Downtime Summary by APS Groups, 08/09/2004 to 08/16/2004:

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

446

08/09/04

05-BM

Closed

35 min

ASD-SI

PSS Minor Fault #475:C_DR2_CLSE_FAIL. User requested that ASD-SI look at the door between scans.

Air cylinder required adjustment to air pressure. CCWP was closed and the beamline returned to operation.

448

08/15/04

02-BM

Open

2 hr 30 min

AOD-IT

User at 2-BM having trouble with computers. Computer running Windows cannot connect to another computer running a cluster where data is saved.

J. Stevens and B. Sheehan contacted C. Roehrig and worked out the problems. Chris said the computer stations were working properly right at 5 p.m.

Users can now view their own downtime by accessing their sector specific records:   http://beam.aps.anl.gov/pls/apsweb/wam0002.efo_main_menu

Search Downtime Records tool will allow Users to view records for the beamline in which they are associated.

Safety:   E. Chang reminds the group of upcoming safety facility inspections in 434 & 435 beginning August 18, 2004.   If you have any requests, please contact your floor coordinators.

B. Glagola notes that ESAFs are still not being turned in with enough time for all approvals before beginning.   He will circulate a memo to safety coordinators outlining policies and procedures for posting ESAFs.  

Maintenance Period R. Salazar:   The information concerning maintenance activities below will also go out on cat-net email.

Work affecting beamlines during the August/September Maintenance Period (8/27/04 through 9/27/04 )   Posted 8/16/04

Switch Gear Maintenance X-7 - Saturday August 28

Sectors 24 , 25, 26 & 27

Regular electrical power will be off from approximately 7 a.m. to 6 p.m.  

FC will contact beamline personnel and establish some 120 V power to their area for necessary items.

Emergency lights and power will be available except for two short durations during transfers (<15 seconds).

Transfer Switch Testing: Saturday August 28

Two short E-power transfer interruptions (<15 seconds) following the schedule below:  

Starting at 6:15 a.m. and approximately every 20 minutes to do the next transfer switch.

X-4        LOM 434, sectors 12, 13, 14 and 15

X-5        LOM 435, sectors 16, 17, 18 and 19

X-8        Sectors 28, 29, 30 and 31

At ~7:30 a.m. power to X-7 switchgear will be shutdown.

X-9        LOM 438, sectors 32, 33, 34 and 35

A-1, R-1 (Bldg 420), J-1 (Bldg 412)

X-1        LOM 431 Sectors 1, 2 and 3

X-2        LOM 432, Sectors 4, 5, 6 and 7

X-3        LOM 433, Sectors 8, 9, 10 and 11

At ~11:00 a.m., Bldg 450 will be shut down ( no process water or chilled water flow to whole APS site ) and the transfer switch will be tested.   Normal 450 operations should resume at ~1:00 p.m.

X-7 power will be restored between 5:30 p.m. and 6:00 p.m.

Switch Gear Maintenance X-6 - Saturday September 11

Sectors 20 , 21, 22 & 23 and LOM 436

Regular electrical power will be off from approximately 6 a.m. to 6 p.m.  

FC will contact beamline personnel and establish some 120 V power to their area for necessary items.

Emergency lights and power will be available except for two short durations during transfers (<15 seconds).

Front End

33-BM:   Replace beryllium window assembly and bake out, need to bleed water system & remove water connectors.   WR# 15608

7-ID:   Replace commissioning window assembly, FEV and bake out second half of front end from SV, need to bleed water system and disconnect X-ray BPM connectors.   WR# 15607

17-BM, 17-ID , 18-BM & 18-ID:   Bake out 1/2 of the front end at a temperature of 130C (from the new exit valve installed in the storage ring to the slow valve).   WR# 15597

19-BM, 19-ID, 20-BM & 20-ID:   Bake out 1/2 of the front end at a temperature of 130C (from the new exit valve installed in the storage ring to the slow valve).   WR# 15596

23-BM & 30-ID:   Install new pneumatics.   WR# 15603 & WR# 15604

23-BM:   Installation of FE components.   WR# 15428

24-BM:   Remove wall plug, install wall collimator tube, align wall collimator tube, and install lead shielding around wall collimator tube.   WR# 15427

30-ID:   Install front end at sector 30 ID, bake out, align, and complete EPS/PSS required to validate front end.

Survey and Alignment

29-ID and 30-ID:   Implement the Decker distortion.   WR# 15494

Provide S&A support for FE installations.

ID Work

3-ID (3US - Remove, re-measure and reinstall ID)

Calibration and maintenance of all IDs

Water

7-ID:   Connect beamline water to APS supplied process water.   WR# 15635

12-ID, 13-ID and 20-ID:   Replace Viatran flow transmitters in beamline hutches in12-ID, 13-ID and 20-ID with new Yokagawa flow transmitters to be located exterior to hutch.   This is part of reliability upgrade.   Water piping/hoses to components will not be removed for this work. WR# 15619

16-ID:   Connect chilled water heat exchanger to APS supplied process water.   WR# 15638

20-ID:   Connect beamline water to APS supplied process water.   WR# 15636

PSS

Re-Validations

Dates

Beamlines

August 30-31

2-ID

August 30-31

4-ID

August 30-31

5-BM

September 7-8

6-ID

September 7-8

8-ID

September 9-10

10-ID

September 13-14

12-ID

September 13-14

34-ID

September 16-17

35-BM

September 16-17

35-ID

2-ID, 4-ID, 5-BM, 6-ID, 8-ID, 10-ID, 12-ID, 34-ID, 35-BM & 35-ID.   Replace watchdog timer relays with solid-state relay equivalents.   WR# 15588

10-ID.   Install remote shutter interface and associated wiring.   WR# 15589

If you require cryo -pump maintenance, and/or enclosure door maintenance during this period, please contact your Floor Coordinator to make arrangements.

TWG:   There will be a TWG meeting on Thursday, August 19, beginning 10:00 a.m.   This meeting will take place in conference room A1100.   The agenda follows:

Facility Update and News:   R.Klaffky , AOD
Presentations and Discussions:  
Time-resolved experiments at Sector-5 - J.Quintana , DND-CAT
Update on the TWG Subgroup for technology development in macromolecular crystallography - M.Capel , NE-CAT
Framework integrating Blu -Ice with EPICS - A.Urakchin , GM/CA-CAT